OpenStudy (anonymous):

PARABOLA QUESTION! Write the equation in standard form Y=16x^2+32x+5

OpenStudy (ipwnbunnies):

:o Looks like it is.

OpenStudy (anonymous):

No, no parabola standard form

OpenStudy (ipwnbunnies):

That is. y = Ax^2 + Bx + C is standard form for a quadratic equation.

OpenStudy (anonymous):

as in (x-h)^2=4p(y-k). I didn't ask for quadratic form.

OpenStudy (ipwnbunnies):

Facepalm. That's not standard form.

OpenStudy (anonymous):

If this helps, I have the answer: (x-1)^2=-1/16(y-21)

OpenStudy (anonymous):

AWMIGAWD! I know what quadratic form is! I took that 3 years ago! This is "Standard form of Equations for parabolas".

OpenStudy (anonymous):

O.o

OpenStudy (anonymous):

@ranga so how did they get (x-1)^2=-1/16(y-21) as an answer for this?

OpenStudy (ranga):

That is not the correct answer to this problem. They are two different parabolas. There is some problem with this question. y = 16x^2 + 32x + 5 is already in standard form and they ask you to put it in standard form. Then you show the answer which is a different parabola than the one given. You can cross-multiply and simplify the answer and you won't get the original parabola. I will try to put it in the vertex form by completing the square: y = 16x^2 + 32x + 5 y = 16(x^2 + 2x + 5/16) y = 16{ (x + 1)^2 - 1 + 5/16 } y = 16{ (x+1)^2 - 11/16 } y = 16(x+1)^2 - 11 This parabola has vertex at (-1, -11). It is an upright parabola that opens upward. (The answer is an upside down parabola that opens downward.)

OpenStudy (anonymous):

Gee, Thanks a lot man! How'd you complete the square so simply? I usually take up a whole page completing the square! @ranga

OpenStudy (ranga):

To complete the square of say, x^2 + 6x - 8 You divide the coefficient of x by 2: 6/2 = 3. This 3 will go within the square: (x + 3)^2 Then you subtract 3^2 So x^2 + 6x - 8 = (x + 3)^2 - 3^2 - 8 = (x+3)^2 - 17

OpenStudy (ranga):

x^2 - 14x + 8 = (x - 7)^2 - 7^2 + 8 = (x-7)^2 - 49 + 8 = (x - 7)^2 - 41

OpenStudy (anonymous):

Don't you have to add the (b/2)^2 to both sides as well?

OpenStudy (ranga):

If you have an equation: x^2 - 14x + 8 = 0 then you add (b/2)^2 to both sides. But when you have just an expression (there are no two sides since there is no equal sign) then you subtract it.

OpenStudy (ranga):

Adding to the right is same as subtracting from the left.

OpenStudy (anonymous):

Wait. I never fully understood this, I always asked my teacher and she would get pissed off, hopefully you don't see it as a stupid question too haha. How come, when you add and subtract b/2^2 from the same side it doesn't cancel out? :o

OpenStudy (ranga):

Example: Complete the square of x^2 + 14x - 3 Divide coefficient of x by 2: 14/2 = +7 This +7 will go inside the square: (x+7)^2. To understand why we subtract 7^2, expand (x+7)^2 and see what you get: (x+7)^2 = x^2 + 14x + 7^2. Compare this to what we started with: x^2 + 14x - 3. By squaring (x+7)^2, we have introduced an extra term 7^2 in the expansion of (x+7)^2. So we need to subtract 7^2. x^2 + 14x - 3 = (x+7)^2 - 7^2 - 3 = (x+7)^2 - 49 - 3 = (x+7)^2 - 52

OpenStudy (anonymous):

OHHHHHH! Wow. Thank you so much ranga. I really appreciate the time you spent helping me out.

OpenStudy (ranga):

You are welcome.